# Cohort-Individual Cooperative Learning for Multimodal Cancer Survival Analysis

### v1

@@ 3.2 Cohort-individual Cooperative Learning
3.3 Multimodal Fusion and Prediction
Recently, Transformer [42] has shown impressive performance on multimodal learning, and thus is employed to integrate the decomposed components for survival predictions. Specifically, we concatenate a class token $U$ with the decomposed components $G$, $P$, $C$, and $S$ as the inputs of Transformer $\Phi_{t}$. Moreover, a fully-connected layer with Sigmoid activation, denoted as $\sigma$, is appended to the output of class token for predicting the hazard function $H$:
$H=\sigma(\Phi_{t}(U,G,P,S,C)).$ (7)
For survival prediction, following previous works [25, 26, 28, 29], we construct an one-hot time-series vector $T$ based on $n$ equal parts of ground truth survival time. Under this setting, the original event time regression problem is simplified to a classification problem. We find out the time interval $t_{k}$ in which the ground truth event occurred as the classification label $k$ for each patient. Each patient sample is defined as a triplet $\{H,c,k\}$, where $H=\{h_{1},...h_{n}\}$ is the predicted hazard vector measuring the probability of ground truth event time located in corresponding time interval $t_{k}$. Additionally, we define the discrete survival function $f_{sur}(H,k)=\prod_{j=1}^{k}(1-h_{j})$. Following previous works [28, 29, 25], we generalizes the negative log-likelihood (NLL) with censorship to supervise the survival prediction by:
Finally, the overall loss function of our framework is:
$L=L_{surv}+\alpha L_{cohort},$ (9)
where $\alpha$ controls the effect of our cohort guidance.
@@ 4.1 Experiment Setups
Dataset. To validate the effectiveness of the proposed methods, we evaluate our framework on five datasets from TCGA: Bladder Urothelial Carcinoma (BLCA, n=372), Breast Invasive Carcinoma (BRCA, n=956), Glioblastoma & Lower Grade Glioma (GBMLGG, n=569), Lung Adenocarcinoma (LUAD, n=453) and Uterine Corpus Endometrial Carcinoma (UCEC, n=480). These datasets contain hundreds of paired genomics, pathological images, and follow-up data collected from multiple centers, enabling center-agnostic analysis on precision oncology. We collected all diagnostic WSIs used for primary diagnosis, resulting in 2,830 WSIs with an average of 15k patches per WSI at 20x magnification (assuming 256 x 256 patches). Genomic profiles provide molecular information of individual, including RNA sequencing (RNA-seq), Copy Number Variation (CNV), Simple Nucleotide Variation (SNV), DNA methylation. Following previous works [25, 29], we use RNA-seq, CNV, and SNV sequences and further group them into six genomic sub-sequences: 1) Tumor Suppression, 2) Oncogenesis, 3) Protein Kinases, 4) Cellular Differentiation, 5) Transcription, and 6) Cytokines and Growth.
Evaluation Metrics. The concordance index (C-index) is the proportion of all comparable patient pairs for which the predicted outcome is consistent with the actual outcome. In survival analysis, the predicted outcome is considered consistent with the actual outcome if the predicted survival time is longer for individuals who indeed have a longer survival time, compared to the predicted event time for others. Moreover, Kaplan-Meier and T-test analyses are employed to assess the significance of differences in survival predictions between high- and low-risk groups.
Implementation. For each dataset, we adopted five-fold cross-validation to evaluate our model and other compared methods. Our framework was implemented using PyTorch running on a single NVIDIA GTX 3090 GPU. For optimization, we employed the SGD optimizer with a learning rate of 1e-3 to train our framework for 30 epochs.
@@
Figure 6: Kaplan-Meier curves of the proposed framework on five cancer datasets.
@@
Figure 7: T-test analysis of the proposed framework on five cancer datasets.
4.2 Comparative Methods
  • Porpoise [27] is a multimodal fusion model that aggregates WSI patch embeddings with the generated attention scores and processes genomic profiles using a SNN network. The extracted multimodal features are simply concatenated to produce survival predictions.
  • SurvPath [26] learns biological pathway tokens from transcriptomics that can encode specific cellular functions and fuse two modalities using a memory-efficient multimodal Transformer to model interactions between pathway and histology patch tokens.
@@ 4.2 Comparative Methods
  • CMTA [29] explores the intrinsic cross-modal correlations and transfer potential complementary information by enhancing modality-specific representations by integrating with multimodal representations.
In addition, we build several models for a comprehensive comparison, including: 1) SNNTrans: two SNN layers followed with a transformer; 2) MaxMIL/MeanMIL: a maximum/average operation to aggregate multiple instances; and 3) DualTrans: combining SNN and TransMIL for genomic profile and pathological images, respectively, while a Transformer is employed for multimodal fusion. We unify the settings of compared models to make a fair comparison and report the re-produced results in our own environment.
@@
4.3 Quantitative Evaluation
C-index Comparison. As shown in Tab. 1, the proposed framework achieves an average score of 72.60%, which surpasses existing methods by a significant margin. On each dataset, our model obtains 68.62%, 68.40%, 86.14%, 69.57%, and 70.26% C-index scores on BLCA, BRCA, GBMLGG, LUAD, and UCEC, which are state-of-the-art performance on most datasets, while the second best on BLCA dataset with a small margin to the best 69.10% reported by CMTA [29]. Compared to unimodal models, either genomics- or pathology-based, our framework achieves significantly improvements on all datasets. It indicates that our framework can effective integrate complementary information in genomic profiles and pathology images. Compared to multimodal models, our framework reports more robust and significant performance on survival analysis. For example, CMTA [29] performs well on the BLCA dataset but still inferior to our method on others.
@@ 4.3 Quantitative Evaluation
Models Datasets Overall
CCA MKD CGM BLCA BRCA GBMLGG LUAD UCEC
(Baseline) 0.6607±0.0319 0.6637±0.0621 0.8393±0.0174 0.6706±0.0343 0.6724±0.0192 0.7013
0.6701±0.0287 0.6767±0.0532 0.8433±0.0241 0.6792±0.0417 0.6844±0.0331 0.7107
0.6783±0.0296 0.6718±0.0416 0.8511±0.0233 0.6855±0.0397 0.6872±0.0388 0.7147
0.6862±0.0253 0.6840±0.0339 0.8614±0.0149 0.6957±0.0231 0.7026±0.0475 0.7260
Kaplan-Meier Analysis. The Kaplan-Meier (KM) analysis is a non-parametric statistic used to estimate the survival function from lifetime data. Specifically, we use the median survival time of the entire cohort to divide all patients into high- and low-risk groups, represented as red and blue curves, respectively. The p-value in KM analysis indicates the probability of observing the observed difference in survival rates between two groups, under the null hypothesis that there is no true difference between them. As shown in Fig. 6, the p-values achieved by our framework are significantly lower than 0.05 on all five datasets, which indicates a statistically significant discrimination between high- and low-risk groups. We also compare the p-values with several representative methods in Tab. 2. Our model achieves the lowest p-values on BLCA and BRCA, meanwhile is competitive to state-of-the-art methods on GBMLGG, LUAD, and UCEC. Our framework can generalize across different cancers to produce robust predictions, holding the potential to significantly advance both clinical practice and cancer research.
T-test Analysis. The T-test is a statistical analysis method used to determine if there is a significant difference between the means of two groups. In the T-test analysis, the t-value quantifies the size of the difference relative to the variability within two groups, while the p-value is the probability that the results from our results occurred by chance. As the results shown in 7, our framework can distinguish between different groups with good p-values and t-values, especially on BLCA, BRCA, and GBMLGG datasets (p-value <  0.05). However, LUAD and UCEC datasets are still challenging for our framework (p-value >  0.05).
Figure 8: T-SNE visualization of the knowledge distribution.
@@ 4.4 Ablation study
Module Analysis. To demonstrate the effectiveness of our methods, we progressively append our modules to our baseline DualTrans, including Cluster Center Alignment (CCA), Multimodal Knowledge Discrimination (MKD), and Cohort Guidance Modeling (CGM). As shown in Tab. 3, the proposed methods effectively and consistently improve the performance of the baseline DualTrans model. CCA aligns the cluster centers after K-means, facilitating the extraction of crucial information from important patches. Moreover, MKD enables more effective multimodal fusion by comprehensively decomposing multimodal knowledge, reducing redundancy and reinforcing valuable information. Further, CGM leverages the cohort guidance to promote a general multimodal interaction learning. Collaborating these modules, our full framework achieves significant improvements over the baseline model.
@@
Knowledge Component Visualization. In Fig. 8, we visualize the feature distributions after MKD using T-SNE [43], where the orange, green, blue and purple points indicate common, synergistic, genomic-specific, and pathology-specific knowledge, respectively. The distributions of these components differ from each other and are consistent with our intuition that common and synergistic features lie in the middle of two modality-specific features. Moreover, the features of each component are clustered as several groups, which demonstrates that patients of similar risk tend to be closer than other patients under our cohort guidance. For example, genomic-specific features are basically clustered as four groups, which is consistent with the group number used in our framework. It proves that our framework can extract discriminative representations to facilitate multimodal fusion.
@@
Hyper-parameter Experiment. There are several hyper-parameters used in our framework, including $k$ in K-means, anchor updating ratio $\tau$, cohort bank length $b$, and the number of patient groups $r$. We provide experiments results on these hyper-parameters in Tab. 4. From these results, we conclude that $k=6$, $\tau=0.1$, $b=10$ and $r=4$ work best in our framework, and thus are employed as the main settings of our framework. Moreover, smaller values of $k$, $b$, and $r$ imply fewer features and computational loads, but the performance usually drops significantly. When exceeding adequate values, enlarging them may not offer additional improvements. For $\tau$, it controls the updating rate of anchor, which is used in cluster center alignment. After a period of training, the anchor becomes stable, so the effect of $\tau$ may not be significant.
Knowledge Attention Analysis. To further illustrate the differences between knowledge components, we visualize their attention on input data in Fig. 9. In sub-figures (a)-(c), we visualize the attention regions of redundancy, synergy, and pathology-specific knowledge on input pathology images, while the last sub-figure (d) shows attention scores on genomic profiles. For pathology images, the focusing regions vary from component to component, indicating the distinctiveness between them. For genomic profiles, the first sample focus more on synergy, whereas the second one has a stronger reliance on redundancy. It implies that two types of multimodal interactions are necessary for effective survival analysis.

### v2

@@ 3.2 Cohort-individual Cooperative Learning
3.3 Multimodal Fusion and Prediction
Recently, Transformer [46] has shown impressive performance on multimodal learning, and thus is employed to integrate the decomposed components for survival predictions. Specifically, we concatenate a class token $U$ with the decomposed components $G$, $P$, $C$, and $S$ as the inputs of Transformer $\Phi_{t}$. Moreover, a fully-connected layer with Sigmoid activation, denoted as $\sigma$, is appended to the output of the class token for predicting the hazard function $H$:
$H=\sigma(\Phi_{t}(U,G,P,S,C)).$ (7)
For survival prediction, following previous works [29, 30, 32, 33], we construct a one-hot time-series vector $T$ based on $n$ equal parts of ground truth survival time. Under this setting, the original event time regression problem is simplified to a classification problem. We find out the time interval $t_{k}$ in which the ground truth event occurred as the classification label $k$ for each patient. Each patient sample is defined as a triplet $\{H,c,k\}$, where $H=\{h_{1},...h_{n}\}$ is the predicted hazard vector measuring the probability of ground truth event time located in corresponding time interval $t$. Additionally, we define the discrete survival function $f_{sur}(H,k)=\prod_{j=1}^{k}(1-h_{j})$. Following previous works [32, 33, 29], we generalize the negative log-likelihood (NLL) with censorship to supervise the survival prediction by:
Finally, the overall loss function of our framework is:
$L=L_{surv}+\alpha L_{cohort},$ (9)
where $\alpha$ controls the effect of our cohort guidance.
@@ 3.3 Multimodal Fusion and Prediction
Pathology feature extraction backbone – HIPT [52]
MCAT [29] Multimodal 0.6483±0.0218 0.6364±0.0277 0.8559±0.0089 0.6587±0.0254 0.6715±0.0602 0.6941
SurvPath [30] Multimodal 0.6295±0.0295 0.5821±0.0675 0.8474±0.0098 0.6384±0.0326 0.6457±0.0589 0.6686
MOTCat [32] Multimodal 0.6556±0.0292 0.6714±0.0446 0.8514±0.0107 0.6694±0.0371 0.7044±0.0643 0.7104
CMTA [33] Multimodal 0.6542±0.0373 0.6545±0.0784 0.8657±0.0149 0.6872±0.0129 0.6787±0.0449 0.7081
Ours Multimodal 0.6651±0.0251 0.6621±0.0452 0.8667±0.0098 0.6775±0.0226 0.6933±0.0408 0.7124
Pathology feature extraction backbone – PLIP [53]
MCAT [29] Multimodal 0.6549±0.0175 0.6761±0.0456 0.8388±0.0240 0.6651±0.0163 0.7029±0.0519 0.7076
SurvPath [30] Multimodal 0.6382±0.0483 0.6252±0.0317 0.8438±0.0208 0.6599±0.0448 0.6938±0.0747 0.6922
MOTCat [32] Multimodal 0.6628±0.0354 0.6698±0.0664 0.8381±0.0144 0.6796±0.0381 0.6845±0.0480 0.7069
CMTA [33] Multimodal 0.6586±0.0192 0.6834±0.0559 0.8515±0.0271 0.6701±0.0241 0.7300±0.0526 0.7187
Ours Multimodal 0.6633±0.0445 0.6797±0.0695 0.8578±0.0235 0.6784±0.0383 0.7309±0.0478 0.7220
Pathology feature extraction backbone – CONCH [54]
MCAT [29] Multimodal 0.6605±0.0159 0.6372±0.0571 0.8438±0.0117 0.6640±0.0275 0.7075±0.0373 0.7026
SurvPath [30] Multimodal 0.6422±0.0194 0.6740±0.0445 0.8447±0.0237 0.6438±0.0493 0.7477±0.0333 0.7105
MOTCat [32] Multimodal 0.6662±0.0327 0.6526±0.0294 0.8442±0.0188 0.6736±0.0371 0.7160±0.0319 0.7105
CMTA [33] Multimodal 0.6694±0.0148 0.7273±0.0613 0.8518±0.0206 0.6750±0.0100 0.7431±0.0484 0.7333
Ours Multimodal 0.6701±0.0323 0.6958±0.0431 0.8454±0.0186 0.6868±0.0253 0.7649±0.0343 0.7326
Pathology feature extraction backbone – UNI [55]
MCAT [29] Multimodal 0.6510±0.0291 0.6640±0.0425 0.8515±0.0289 0.6566±0.0306 0.7007±0.0556 0.7047
SurvPath [30] Multimodal 0.6290±0.0376 0.7005±0.0521 0.8286±0.0204 0.6321±0.0550 0.7376±0.0475 0.7056
MOTCat [32] Multimodal 0.6525±0.0280 0.6651±0.0421 0.8491±0.0154 0.6684±0.0275 0.7383±0.0552 0.7147
CMTA [33] Multimodal 0.6521±0.0319 0.7059±0.0304 0.8541±0.0207 0.6642±0.0462 0.7403±0.0432 0.7233
Ours Multimodal 0.6404±0.0314 0.7152±0.0410 0.8536±0.0126 0.6737±0.0395 0.7613±0.0482 0.7288
@@ 4.1 Experiment Setups
Dataset. To validate the effectiveness of the proposed methods, we evaluate our framework on five datasets from TCGA: Bladder Urothelial Carcinoma (BLCA, n=372), Breast Invasive Carcinoma (BRCA, n=956), Glioblastoma & Lower Grade Glioma (GBMLGG, n=569), Lung Adenocarcinoma (LUAD, n=453) and Uterine Corpus Endometrial Carcinoma (UCEC, n=480). These datasets contain hundreds of paired genomics, pathological images, and follow-up data collected from multiple centers, enabling center-agnostic analysis on precision oncology. We collected all diagnostic WSIs used for primary diagnosis, resulting in 2,830 WSIs with an average of 15k patches per WSI at 20x magnification (assuming 256 x 256 patches). Genomic profiles provide molecular information about individuals, including RNA sequencing (RNA-seq), Copy Number Variation (CNV), Simple Nucleotide Variation (SNV), and DNA methylation. Following previous works [29, 33], we use RNA-seq, CNV, and SNV sequences and further group them into six sub-sequences: 1) Tumor Suppression, 2) Oncogenesis, 3) Protein Kinases, 4) Cellular Differentiation, 5) Transcription, and 6) Cytokines and Growth.
Evaluation Metrics. The concordance index (C-index) is the proportion of all comparable patient pairs for which the predicted outcome is consistent with the actual outcome. In survival analysis, the predicted outcome is considered consistent with the actual outcome if the predicted survival time is longer for individuals who indeed have a longer survival time, compared to the predicted event time for others. We also calculate the standard deviations of C-index scores across five folds. Moreover, Kaplan-Meier and T-test analyses are employed to assess the significance of differences in survival predictions between high- and low-risk groups.
Implementation. For each dataset, we adopted five-fold cross-validation to evaluate our model and other compared methods. Specifically, first, we shuffle the dataset randomly and split the dataset into five groups. For each group, we take this group as a test data set, while the remaining groups as a training data set. We fit the models on the training set and evaluate them on the test set to report the C-index scores. Our framework was implemented using PyTorch running on a single NVIDIA GTX 3090 GPU. For optimization, we employed the SGD optimizer with a learning rate of 1e-3 to train our framework for 30 epochs. We unify the settings of compared models to make a fair comparison and report the re-produced results in our environment. For the initialization of random seed, we set it to 0 in all experiments.
4.2 Comparative Methods
  • MCAT [29] is a multimodal fusion model that utilizes the co-attention between genomic and pathology features as the attention vector to select informative patches in WSIs. After that, two Transformers are employed to fuse genomic and pathology features, respectively. Finally, two features are concatenated to produce survival predictions.
  • SurvPath [30] learns biological pathway tokens from transcriptomics that can encode specific cellular functions and fuses two modalities using a memory-efficient multimodal Transformer to model interactions between pathway and histology patch tokens.
@@ 4.2 Comparative Methods
    Figure 6: Kaplan-Meier curves of the proposed framework on five TCGA datasets.
@@
    Figure 7: T-test analysis of the proposed framework on five TCGA datasets.
@@
  • CMTA [33] explores the intrinsic cross-modal correlations and transfers potential complementary information by enhancing modality-specific representations by integrating with multimodal representations.
In addition, we build several models for a comprehensive comparison, including 1) SNNTrans: two SNN layers followed by a transformer; 2) MaxMIL/MeanMIL: a maximum/average operation to aggregate multiple instances; and 3) DualTrans: combining SNN and TransMIL for genomic profile and pathological images, respectively, while a Transformer is employed for multimodal fusion.
4.3 Quantitative Evaluation
C-index Comparison. As shown in Tab. 1, the proposed framework achieves an average score of 72.60%, which surpasses existing methods by a significant margin. On each dataset, our model obtains 68.62%, 68.40%, 86.14%, 69.57%, and 70.26% C-index scores on BLCA, BRCA, GBMLGG, LUAD, and UCEC, which are state-of-the-art performance on most datasets, while the second best on BLCA dataset with a small margin to the best 69.10% reported by CMTA [33]. Compared to unimodal models, either genomics- or pathology-based, our framework achieves significant improvements on all datasets. It indicates that our framework can effectively integrate complementary information in genomic profiles and pathology images. Compared to multimodal models, our framework reports a more generalized performance on survival analysis. For example, CMTA [33] performs well on the BLCA dataset but is still inferior to our method on others.
To prove the robustness of our framework, we compare the proposed framework with main competitors, including MCAT [29], SurvPath [30], MOTCat [32], and CMTA [33] by using latest pathology foundation models, including HIPT [52], PLIP [53], UNI [55], and CONCH [54]. Our framework achieves the best overall performance when using HIPT, PLIP, and UNI as pathology feature extractors, respectively. Furthermore, our framework remains competitive to the best performance achieved by CMTA when using CONCH. These experiments further prove the robustness of the proposed framework.
Kaplan-Meier Analysis. The Kaplan-Meier (KM) analysis is a non-parametric statistic used to estimate the survival function from lifetime data. Specifically, we use the median survival time of the entire cohort to divide all patients into high- and low-risk groups, represented as red and blue curves, respectively. The p-value in KM analysis indicates the probability of observing the observed difference in survival rates between two groups, under the null hypothesis that there is no true difference between them. As shown in Fig. 6, the p-values achieved by our framework are significantly lower than 0.05 on all five datasets, which indicates a statistically significant discrimination between high- and low-risk groups. We also compare the p-values with several representative methods in Tab. 2. Our model achieves the lowest p-values on BLCA and BRCA and, meanwhile is competitive to state-of-the-art methods on GBMLGG, LUAD, and UCEC. Our framework can generalize across different cancers to produce robust predictions, holding the potential to significantly advance both clinical practice and cancer research.
@@ 4.3 Quantitative Evaluation
T-test Analysis. The T-test is a statistical analysis method used to determine if there is a significant difference between the means of two groups. In the T-test analysis, the t-value quantifies the size of the difference relative to the variability within two groups, while the p-value is the probability that the results from our results occurred by chance. As the results shown in Fig. 7, our framework can distinguish between different groups with good p-values and t-values, especially on BLCA, BRCA, and GBMLGG datasets (p-value <  0.05). However, LUAD and UCEC datasets are still challenging for our framework (p-value >  0.05).
Efficiency Comparison. Since all compared survival models require extracting pathology patch embeddings beforehand, we exclude the embedding extraction process in our efficiency comparison in Tab. 3. Our framework demands increased GPU resources and computation time compared to previous methods, primarily due to the inclusion of online K-means clustering. However, by employing K-means clustering in advance, the GPU memory consumption of our pure network (denoted with *) is considerably lower than existing methods, while achieving faster processing speeds than state-of-the-art CMTA and MOTCat methods.
Table 3: Computational resource comparison. Training and test time measures the numbers of milliseconds for processing a single sample. * means that K-means is employed to cluster pathology patches beforehand.
Methods GPU memory Training time Test time
MCAT [29] 4.06 13.7 9.6
SurvPath [30] 1.95 8.9 5.3
MOTCat [32] 3.09 39.2 38.3
CMTA [33] 19.70 33.1 28.1
Ours* 0.44 21.6 16.1
Ours 18.9 95.2 81.1
@@
CCA MKD CGM BLCA BRCA GBMLGG LUAD UCEC Overall
(Baseline) 0.6607 0.6637 0.8393 0.6706 0.6724 0.7013
0.6701 0.6767 0.8433 0.6792 0.6844 0.7107
0.6783 0.6718 0.8511 0.6855 0.6872 0.7147
0.6862 0.6840 0.8614 0.6957 0.7026 0.7260
@@ 4.4 Ablation study
Module Analysis. To demonstrate the effectiveness of our methods, we progressively append our modules to our baseline DualTrans, including Cluster Center Alignment (CCA), Multimodal Knowledge Discrimination (MKD), and Cohort Guidance Modeling (CGM). As shown in Tab. 4, the proposed methods effectively and consistently improve the performance of the baseline DualTrans model. CCA aligns the cluster centers after K-means, facilitating the extraction of crucial information from important patches. Moreover, MKD enables more effective multimodal fusion by comprehensively decomposing multimodal knowledge, reducing redundancy, and reinforcing valuable information. Further, CGM leverages the cohort guidance to promote general multimodal interaction learning. Collaborating with these modules, our full framework achieves significant improvements over the baseline model.
(a) BLCA
(b) BRCA
(c) GBMLGG
Refer to captionRefer to caption
(d) LUAD
(e) UCEC
Figure 8: T-SNE visualization of the knowledge distributions.
Impact of Cluster Alignment. To prove the necessity of cluster alignment, we calculate the mean L2 distances among cluster representations at the same ordinal position before and after our alignment process in Tab. 5. Our alignment method significantly reduces the distances between cluster representations. However, there still exist considerable distances between these representations, mainly due to the different patch distributions and personalized information of patients. Addressing this issue may necessitate the development of more sophisticated approaches in future research.
Impact of Knowledge Encoders. In our framework, we employ two encoders to learn common and synergistic knowledge, respectively. As introduced in previous sections and many related works [42, 43, 44, 45], the knowledge in multimodal fusion may be one of common, synergistic, and two modality-specific components, serving as the theoretical foundation of our framework structure. If more co-attention encoders are introduced, they will not fall out of these four knowledge components, becoming redundant information. To prove this point, we provide the results of our framework with more co-attention encoders in Tab. 6, where two of them are supervised as common and synergistic encoders and others have no specific supervision signals. Furthermore, we also validate our framework with removed synergistic or common encoders. With more encoders employed, we have not seen substantial improvements in overall performance. Meanwhile, since the integrity of multimodal knowledge is destroyed, we witness varying performance degradation on different datasets.
Table 5: Average distances of cluster representations before and after alignment.
Distances BLCA BRCA GBMLGG LUAD UCEC
Before alignment 4.6501 5.5141 7.1205 5.6450 4.6632
After alignment 3.4775 4.6234 6.0795 4.9767 3.8678
Table 6: Ablation study on our framework with different numbers of knowledge encoders.
# of encoders BLCA BRCA GBMLGG LUAD UCEC Overall
5 (Ours + 3) 0.6689 0.6931 0.8513 0.6856 0.7089 0.7216
3 (Ours + 1) 0.6883 0.6798 0.8581 0.6967 0.6998 0.7245
1 (Common) 0.6793 0.6851 0.8549 0.6772 0.6967 0.7186
1 (Synergistic) 0.6764 0.6489 0.8576 0.6752 0.6794 0.7075
2 (Ours) 0.6862 0.6840 0.8614 0.6957 0.7026 0.7260
Knowledge Component Visualization. In Fig. 8, we visualize the feature distributions after MKD using T-SNE [56], where the orange, green, blue, and purple points indicate common, synergistic, genomic-specific, and pathology-specific knowledge, respectively. The distributions of these components differ from each other and are consistent with our intuition that common and synergistic features lie in the middle of two modality-specific features. Moreover, the features of each component are clustered as several groups, which demonstrates that patients of similar risk tend to be closer than other patients under our cohort guidance. For example, genomic-specific features are clustered into four groups, which is consistent with the group number used in our framework. It proves that our framework can extract discriminative representations to facilitate multimodal fusion.
Impact of $l_{k}$. An ablation study on the similarity constraints in $l\_k$ is conducted by removing the supervision signals for different encoders, respectively, as shown in Tab. 7. Since different tasks may benefit more from different knowledge components, removing the similarity constraints causes performance degradation to different extents. It proves that, without explicit modeling of knowledge components, important knowledge in multimodal data is easy to be overlooked.
Table 7: Ablation study on our framework with different similarity constraints in $l_{k}$. G, P, C, and S indicate the constraints for corresponding knowledge components.
Similarity constraints BLCA BRCA GBMLGG LUAD UCEC Overall
G+P+S 0.6598 0.6931 0.8647 0.6722 0.6793 0.7138
G+P+C 0.6784 0.6872 0.8657 0.6868 0.6983 0.7233
P+C+S 0.6720 0.6970 0.8489 0.7003 0.6781 0.7193
G+C+S 0.6915 0.6786 0.8586 0.6834 0.7057 0.7236
G+P+C+S (Ours) 0.6862 0.6840 0.8614 0.6957 0.7026 0.7260
Hyper-parameter Experiment. There are several hyper-parameters used in our framework, including $k$ in K-means, anchor updating ratio $\tau$, cohort bank length $b$, and the number of patient groups $r$. We provide experiment results on these hyper-parameters in Tab. 8. From these results, we conclude that $k=6$, $\tau=0.1$, $b=10$, and $r=4$ work best in our framework, and thus are employed as the main settings of our framework. Moreover, smaller values of $k$, $b$, and $r$ imply fewer features and computational loads, but the performance usually drops significantly. When exceeding adequate values, enlarging them may not offer additional improvements. For $\tau$, it controls the updating rate of the anchor, which is used in cluster center alignment. After a period of training, the anchor becomes stable, so the effect of $\tau$ may not be significant.
@@
Knowledge Attention Analysis. To further illustrate the differences between knowledge components, we visualize their attention on input data in Fig. 9. In sub-figures (a)-(c), we visualize the attention regions of redundancy, synergy, and pathology-specific knowledge on input pathology images, while the last sub-figure (d) shows attention scores on genomic profiles. For pathology images, the focusing regions vary from component to component, indicating the distinctiveness between them. For genomic profiles, the first sample focuses more on synergy, whereas the second one has a stronger reliance on redundancy. It implies that two types of multimodal interactions are necessary for effective survival analysis.